Grade 9Algebra

A vehicle of mass m is moving on a rough horizontal road with momentum P. If the coefficient of friction between the tyre and road is µ, then the stopping distance is?

RO
9 Years agoGrade 9

1 Answer

Darsh Mehta
ApprovedApproved Tutor Answer9 Years ago
Is it( P^2)/m^2 *2ug as velocity=P/m and umg=ma Putting values in v^2=u^2-2as and v=0 we get the above answer